OpenStudy (anonymous):

how to covert 0xAB00FF12 into 0xABFF0012 ? I tried using XOR ...means 0xAB00FF12 ^ 00FFFF00, is this right

OpenStudy (anonymous):

Seems to be just fine. a xor with 0 will produce the same value (0^0=0, 0^1=1), so using the zeros for the parts you want to keep (AB and 12) will work fine. Xorring with a 1 will invert the bits. (1^0=1,1^1=0), so it will indeed flip the middle part 00FF.

OpenStudy (anonymous):

It is formally called binary masking. But you are using hexadecimal and therefore is hard to determine the bitwise action.but if you know your zeros and ones in 0x then your all set.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Mask_(computing)